Table 1            Estimated Total Returns                                                                                                 
State of the Economy Probability T-Bond SETX Golden S&P 500
Recession 5% 5% -19% 20% -14%
Below Average 15% 5% 2% 13% 3%
Average 45% 5% 9% 10% 11%
Above Average 25% 5% 34% 5% 22%
Boom 10% 5% 25% -5% 33%

Calculate the standard deviations and coefficients of variation of returns for the four alternatives. What type of risk do these statistics measure? Is the standard deviation or the coefficient of variation the better measure? How do the alternatives compare when risk is considered?
Answer #1

We need to calculate the mean and standard deviation using the following formula:

Mean of a probability distribution = E(X) = probability * X

Standard deviation = square root ( probability * [X-E(X)]2 )

Coefficient of variation = mean / standard deviation

For T-Bond, the mean = 5% and standard deviation = 0 because the returns do not change. There is no variation because of constant returns, hence coefficient of variation is undefined and this can be confirmed with the formula too. which is 5% / 0 = undefined.

Calculate the mean and standard deviation of SETX, Golden and S&P500 as follows:

E (SETX) = 0.05 * -19 + 0.15 * 2 + 0.45 * 9 + 0.25 * 34 + 0.10 * 0.25 = 14.40%

E (Golden) = 0.05 * -20 + 0.15 * 13 + 0.45 * 10 + 0.25 * 5 + 0.10 * -5 = 8.20%

E (S&P500) = 0.05 * -14 + 0.15 * 3 + 0.45 *11 + 0.25 * 22 + 0.10 * 0.33 = 13.50%

Standard deviation = square root ( probability * [X-E(X)]2 )

Substituting the values for SETX, Golden and S&P500 in the formula above, we get the following values for standard deviation:

Standard deviation (SETX) = 14.12%

Standard deviation (Golden) = 5.64%

Standard deviation (S&P500) = 10.64%

Coefficient of variation (SETX) = Mean / Standard deviation = 14.40 / 14.12 = 0.9802

Coefficient of variation (Golden) = Mean / Standard deviation = 8.20 / 5.64 = 0.6883

Coefficient of variation (S&P500) = Mean / Standard deviation = 13.50 / 10.64 = 0.7883

What type of risk do these statistics measure?

Both Standard deviation and coefficient of variation measures the variability of the data set or dispersion of data around the mean.

Is the standard deviation or the coefficient of variation the better measure?

The coefficient of variation is a relative measure as it encapsulates both mean and standard deviation whereas standard deviation is an absolute measure. When one has to compare two or more distributions then co-efficient of variation is a better measure than standard deviation because an absolute measure comparison is meaningless.

How do the alternatives compare when risk is considered?

SETX is most risky with standard deviation of 14.12%, whereas TBond is risk free. Golden ranks 2nd in risk and S&P500 is 3rd in risk.